Scores of an IQ test have a​ bell-shaped distribution with a mean of 100 and a standard deviation of 17 . Use the empirical rule to determine the following. ​

(a) What percentage of people has an IQ score between 66 and 134 ​?

​(b) What percentage of people has an IQ score less than 49 or greater than 151 ​?

​(c) What percentage of people has an IQ score greater than 151 ​?
Answer #1

a)

P (66<X134) 0.9544 Explanation Step 1: Sketch the curve. The probability that 66 < X 134 is equal to the blue area under the

b)

P = 1-P(49<X<151)

P (49 <X151) 0.9974 Explanation Step 1: Sketch the curve. The probability that 49 <X151 is equal to the blue area under the c

Prob = 1-0.9974 = 0.0026

c)

P (X > 151) 0.0013 Explanation Step 1: Sketch the curve. The probability that X > 151 is equal to the blue area under the cur